A 41 kg skateboarder running at 4.5 m/s [N] jumps onto a 4.4 kg skateboard moving at 2.5 m/s [N]. How fast will the skateboarder and board be traveling once she jumps on the board?

Your Answer:


1
Expert's answer
2021-10-28T08:54:21-0400

We can find the velocity of skateboarder and board from the law of conservation of momentum:


"m_1v_1+m_2v_2=(m_1+m_2)v,""v=\dfrac{m_1v_1+m_2v_2}{m_1+m_2},""v=\dfrac{41\ kg\times4.5\ \dfrac{m}{s}+4.4\ kg\times2.5\ \dfrac{m}{s}}{41\ kg+4.4\ kg}=4.3\ \dfrac{m}{s}."